How do I find equations for lines parallel or perpendicular to y = ⅔ x + 2; (6, -3)
Answer by Boreal(15235)

parallel has same slope
use point slope formula y-y1=m(x-x1), m slope, (x1,y1) point
y+3=(2/3)(x-6), watch the sign
y=(2/3(x)-7
------------------
perpendicular lines have the negative reciprocal slope, so negative reciprocal os 2/3 is -3/2
y+3=(-3/2)(x-6)
y+3=-(3/2)x+9
y=(-3/2)x-12
-----------------

The line going through (0,2) is the line given.